OpenStudy (anonymous):

A lemonade recipe calls for 5 lemons for every 3 cups of water. Find the missing number in the table.  10 11 15 18

OpenStudy (anonymous):

wheres the table

OpenStudy (anonymous):

@Minecraftdude1524 what is the table. You cant answer without it!

OpenStudy (anonymous):

its c based on that its the only multiple of 3 and five

OpenStudy (anonymous):

it is c

OpenStudy (texaschic101):

look at your table...notice that these numbers are equivalent fractions... 5/3 = 15/9 = 20/12

OpenStudy (just_one_last_goodbye):

I would think it's 15 because we multiply 3 * 5 = 15